OpenStudy (kingstone):

Solve for R, please help! 3000 = 2(pi)r50 + (pi)r^2

OpenStudy (hba):

\[3000=100 \pi r+\pi r^2,Right ?\]

OpenStudy (kingstone):

right!

OpenStudy (hba):

\[\pi r^2+100 \pi r -3000=0\] Where, \[a=\pi\] \[b=100 \pi \] c=-3000 in Quad formula.

OpenStudy (kingstone):

This is just and end to a question I was solving, and now I need to find the radius which is r. I plugged in the numbers and solved the other stuff.

OpenStudy (hba):

\[r=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (hba):

I know you can do the rest :) Best of luck and Happy new year buddy :)

OpenStudy (kingstone):

\[r=\frac{ -100\pi \pm \sqrt{100\pi^2-4(-3000\pi)} }{ 2\pi }\] Is that how I should present my answer exactly? By the way, thank you so much, you've already done so much :)

OpenStudy (hba):

\[r=\frac{ -100\pi \pm \sqrt{(100\pi)^2-4(-3000\pi)} }{ 2\pi }\] Now solve further :)

OpenStudy (kingstone):

r=8.78 cm ?

OpenStudy (anonymous):

wait which pi 3.14 or \[\frac{ 22 }{ 7 }\]

OpenStudy (anonymous):

do u already have the answer right

OpenStudy (anonymous):

@karmel its worth noting 22/7 is 3.14 lol

OpenStudy (anonymous):

yea but sometimes it makes a difference whether u put the fraction or the decimal

OpenStudy (anonymous):

if u carry it out further cause the decimal form is 3.141592653589793238462643383279502884197169399375105820974944592307816406286208998348253421170679 where as 22/7 is 3.142

OpenStudy (anonymous):

but they dont give u all of that listen i have my own thing to do and i cant argue right now i have a 90 page packet and i am not even half way done so ok bye bye

OpenStudy (anonymous):

wasnt arguing just joking and alright have fun lmao

OpenStudy (anonymous):

wow really

OpenStudy (anonymous):

?

OpenStudy (anonymous):

nuthin goodbye

OpenStudy (hba):

@Karmel Well you should refer to this if you want to learn about pi http://openstudy.com/users/hba#/updates/50e027c8e4b050087cd0c732

OpenStudy (anonymous):

well i am over this spo no need

OpenStudy (anonymous):

but thank you for your concern

OpenStudy (hba):

Then you would be saying that the value of pi is 355/113,Which infact is the best approximation :) 22/7 is also a good approximation though :)

OpenStudy (anonymous):

ok

OpenStudy (hba):

You're welcome :)